the length of a rectangle is 3 inches less than twice the width. if the perimeter of a recatngle is 18 inches, what is the width of the rectangle
Answer by stanbon(75887)

the length of a rectangle is 3 inches less than twice the width.
---
L = 2W-3
----
if the perimeter of a rectangle is 18 inches, what is the width of the rectangle
Equation:
P = 2(L + W)
18 = 2(2W-3 + W)
9 = 3W-3
3W = 12
Width = 4 inches
Length = 2(4)-3 = 5 inches
===============================
Cheers,
Stan H.
==========